A perchloric acid solution is titrated using barium hydroxide. If the
complete neutralization of 50.0 mL of the HClO4 solution requires 37.0 mL of
0.125 M Ba(OH) 2 , what is the molarity of the HClO4 ?
1
Expert's answer
2019-11-27T07:12:44-0500

Solution.

2HClO4 + Ba(OH)2 = Ba(ClO4)2 + 2H2O

n(Ba(OH)2) = 0.0046 moles

n(HClO4) = 0.0092 moles

C = 0.0092/0.05 = 0.185 M (=0.2 M)

Answer. 0.185 M